Translate the following problems into algebraic equations :

1)Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?

2)A girl buys three apples and seven oranges for $4.56. If an orange costs $0.28 more than an apple, how much does each fruit cost?

Answer by scott8148(6628)

let x=longest side

"The shortest side is 30 yards longer than one-half of the longest side"
__ (x/2)+30

"The second longest side is 2 yards shorter than the longest side"
__ x-2

"The perimeter fence that surrounds the land measures 378 yards"
__ x+(x/2)+30+x-2=378



let x=apple

"an orange costs $0.28 more than an apple" __ x+.28

" three apples and seven oranges for $4.56" __ 3x+7(x+.28)=4.56 __ 3x+7x+1.96=4.56 __ 10x+1.96=4.56